Consider the titration of 0.100 L of 0.200 M ascorbic acid with 0.500 M NaOH. What is the pH at the endpoint of the titration? Enter your answer to two decimal places; for this question we can assume the 5% assumption is valid.
Answer #1

This is a case of titration of a weak acid with a strong base. Let us represent the reaction going on here as

HA + OH- --------> A- + H2O

Let us find out the amount of ascorbic acid present in the solution in mole.

Amount of ascorbic acid = (0.100 L). (0.200 mole/L) = 0.02 mole

NaOH reacts with ascorbic acid on a 1:1 molar ratio; so 0.02 mole of ascorbic acid will be neutralized by 0.02 mole of NaOH.

The volume of NaOH required is (0.02 mole)/(0.500 mole/L) = 0.04 L

At the end point, we have sodium ascorbate and water. We calculate the concentration of sodium ascorbate as (0.02 mole)/(0.14 L) = 0.1428 M (ascorbate is obtained from ascorbic acid on a 1:1 molar ratio).

Now, ascorbate is the conjugate base of a weak acid and will undergo the following equilibrium reaction.

A- + H2O ----------> HA + OH-

Initial                           0.1428                          0       0

Change                         - x                                 x       x

Equilibrium             (0.1428 – x)                   (0+x)(0+x)

The equilibrium constant

Kb = (x).(x)/(0.1428 – x)

Now, the acid dissociation constant for ascorbic acid is Ka = 7.9 x 10-5

We also know Ka.Kb = Kw = 10-14

Hence, Kb = Kw/Ka = (1.0 x 10-14)/(7.9 x 10-5) = 1.2658 x 10-10

Hence, Kb = x2/(0.1428) (We assume x to be very small)

or, x2 = Kb.(0.1428) = (1.2658 x 10-10).(0.1428) = 1.8075 x 10-11

or, x =4.2515 x 10-6

So, the concentration of hydroxide in the solution at the end point is 4.2515 x 10-6.

Now, pOH = - log10[OH-] = -log10(4.2515 x 10-6) = 5.37

Therefore, pH = 14 – pOH = 14 – 5.37 = 8.63

Ans: The pH at the end point is 8.63.